OpenStudy (anonymous):

Write an equation for the table. x -1 0 1 2 y 5 3 1 -1 A. y = -2x + 3 B. y = -x + 3 C. y = 2x + 5 D. y = x + 5

OpenStudy (ajprincess):

Let the equation be y=mx+c Substituting y=5 and x=-1 you get 5=-m+c-(1) substituting y=3 and x=0 you get 3=c (1)=>5=-m+3 5-3=-m -m=2 m=-2 so the eqn is y=-2x+3 Is that clear?

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

yw